# RbRL2.0: Integrated Reward and Policy Learning for Rating-based Reinforcement Learning

### v1

@@ Introduction
Table 1: hyperparameters used in experiments.
(a) HalfCheetah
(b) Walker
(c) Quadruped
Figure 2: Learning curves of different algorithms across three environments. The plots show the mean (solid line) with the standard error (shaded area) over 10 runs.
To evaluate the effectiveness of our proposed method, we compare the new method, labeled as RbRL2.0, with RbRL across 3 DeepMind Control environments (Tassa et al. 2018), namely, HalfCheetah, Walker and Quadruped. These environments are characterized with continuous state and action spaces. Specifically, HalfCheetah is characterized with a 17-dimensional state space, capturing joint positions, velocities, and body orientation, and a 6-dimensional action space, which corresponds to the torques applied to each joint. Walker is characterized with a 24-dimensional state space, representing joint angles, velocities, and torso orientation, and a 6-dimensional action space, controlling forces applied to each limb. Quadruped is characterized with a 78-dimensional state space, including joint positions, velocities, and full body orientation, and a 12-dimensional action space, allowing torque-based control over each leg joint. It is worth noting that HalfCheetah is the simplest environment while Quadruped is the most challenging environment among the three. In other words, we focus on understanding how RbRL2.0 perform in environments with different complex levels. Although there are existing reward functions designed for the three environments, our proposed RbRL2.0 does not require such information. Instead, the original reward functions are used to purely evaluate the performance of the trained policies.
To evaluate the effectiveness of our proposed method, we evaluate our method against RbRL across rating classes of 3, 4, 5, and 6, corresponding to training with 2, 3, 4, and 5 KL divergence terms, respectively. The hyperparameters used in our experiments are provided in Table 1. To ensure reproducibility, each case is implemented with 10 runs using different seeds. The performance of both RbRL and the new RbRL2.0 is evaluated based on the original environment reward. Table 2 presents the average cumulative rewards with standard errors over 10 runs. Figure 2 compares the learning curves of both algorithms across different rating classes. It can be seen from these results that RbRL2.0 consistently outperforms RbRL on Walker and Quadruped while performing similarly as RbRL in HalfCheetah. Meanwhile, RbRL2.0 demonstrates faster learning and achieves higher cumulative rewards in most cases. It is worth mentioning that RbRL2.0 and RbRL perform comparably in the HalfCheetah environment, which indicates the limited value of policy loss in this environment. Moreover, a deeper analysis shows that both RbRL and RbRL2.0 perform best when there are 4 rating classes for the HalfCheetach environment. In other words, the inclusion of more rating classes will adversely impact the performance since the HalfCheetach environment is rather simple. Similarly, the inclusion of new KL loss terms may not help or even adversely impact the performance in RbRL2.0. For more complex environments Walker and Quadruped, RbRL2.0 consistently outperforms RbRL across all rating classes, demonstrating the effectiveness of RbRL2.0 in complex environments where RbRL fails to yield good performance. From Figure 2(c), it can be observed that RbRL2.0 yields a significantly higher cumulative reward than RbRL with a rating class of 5, underscoring its effectiveness in handling more complex environments.
Environment Empirical Return
@@
HalfCheetah $163.02\pm 42.49$ $\mathbf{238.99\pm 34.81}$ $\mathbf{204.59\pm 16.37}$ $\mathbf{235.46\pm 46.10}$
@@
Quadruped $454.39\pm 89.44$ $308.48\pm 62.29$ $227.52\pm 39.92$ $199.83\pm 45.71$
RbRL2.0 (n=3) RbRL2.0 (n=4) RbRL2.0 (n=5) RbRL2.0 (n=6)
HalfCheetah $\mathbf{200.63\pm 45.63}$ $227.13\pm 17.20$ $183.29\pm 15.03$ $205.83\pm 24.63$
@@
Quadruped $\mathbf{420.27\pm 44.00}$ $\mathbf{477.29\pm 69.89}$ $\mathbf{742.05\pm 50.99}$ $\mathbf{306.78\pm 75.69}$
Table 2: Empirical return comparison among different algorithms.
Limitations and Future Work
One limitation of the proposed approach is the variability in individual rating standards, which may be noisy due to differences in how participants interpret and evaluate the environment. Since RbRL relies on users having a basic understanding of the task and environment, these inconsistencies may reduce the reliability of the collected individual ratings. To address this, future work will involve designing a crowdsourcing framework to gather a diverse range of ratings from various participants. Developing effective noise filtering methods is important to ensure the collected data is more representative and robust, improving the reliability and accuracy of the learned models.
Although $\omega_{i}$ should be selected in a descending order with respect to the rating level, it remains an open question how to select the specific value of $\omega_{i}$. Ablation studies will be needed to optimize the performance by selecting the best $\omega_{i}$. Another interesting future research is to integrate the new KL loss terms with other RL methods in dense-reward, sparse-reward, and reward-free settings to further evaluate its effectiveness since the KL loss terms can be implemented with/without rewards. Finally, it is important to develop a systematic approach that uncovers the right number of rating classes for a given environment since a larger number of rating classes may not always lead to better performance. One potential idea is to bring the idea of Just Noticeable Difference, which quantifies the smallest change in a stimulus to be observed by humans, in psychology to determine the boundaries between different ratings. We will explore these directions as part of our future work.

### v2

@@ 5 Experiments and Results
To further evaluate the effectiveness of our proposed method, we compare RbRL-KL against RbRL across rating classes of 3, 4, 5, and 6, corresponding to training with 2, 3, 4, and 5 KL divergence terms, respectively. The hyperparameters used in our experiments are provided in Table 1. To ensure reproducibility, each setting is run 10 times using different random seeds. Table 2 presents the average cumulative rewards with standard errors over 10 runs, and Figure 2 compares the learning curves of both algorithms across different rating classes. The results show that RbRL-KL outperforms RbRL in most environments. In particular, RbRL-KL consistently achieves better performance on Cartpole-balance, Ball-in-cup, HalfCheetah, and Walker, while it only underperforms in Finger-spin and Quadruped when $n=3$. It is worth noting that RbRL outperforms RbRL-KL under the lower rating class because the lower-rated segments form a broad, undifferentiated group. As a result, the KL-divergence terms in RbRL-KL apply a more uniform deviation across all these segments, reducing their impact on policy learning. Therefore, the trained policy is more likely to achieve the local optima. Instead, a more refined rating system will provide better sample separation, allowing the KL-based policy loss in RbRL-KL to learn more effectively. Apart from these specific cases, the proposed RbRL-KL consistently outperforms RbRL, demonstrating its superior performance. To provide a clearer view of the algorithm’s advantages, Table 3 illustrates the percentage improvement achieved by RbRL-KL over conventional reward learning from human ratings. Positive values indicate performance enhancement, showing that RbRL-KL generally improves learning, while minor decreases appear only in Finger-Spin and Quadruped under the low rating class.
Table 2: Empirical return comparison among different algorithms.
Environment ($\uparrow$ complexity) Empirical Return
@@
Cartpole-balance $341.50\pm 47.77$ $402.55\pm 60.65$ $349.84\pm 39.95$ $306.92\pm 39.55$
Ball-in-cup $706.20\pm 104.26$ $789.30\pm 84.62$ $698.15\pm 104.55$ $828.62\pm 57.98$
Finger-spin $\mathbf{530.79\pm 31.15}$ $11.55\pm 24.25 55$$7.88\pm 44.52 55$$59.73\pm 20.49$
HalfCheetah $163.02\pm 42.49$ $238.99\pm 34.81$ $204.59\pm 16.37$ $235.46\pm 46.10$
@@
Quadruped $\mathbf{454.39\pm 89.44}$ $308.48\pm 62.29$ $227.52\pm 39.92$ $199.83\pm 45.71$
RbRL-KL (n=3) RbRL-KL (n=4) RbRL-KL (n=5) RbRL-KL (n=6)
Cartpole-balance $\mathbf{394.62\pm 41.30}$ $\mathbf{417.54\pm 58.89}$ $\mathbf{428.64\pm 43.10}$ $\mathbf{381.79\pm 33.91}$
Ball-in-cup $\mathbf{856.25\pm 77.26}$ $\mathbf{861.47\pm 33.34}$ $\mathbf{790.94\pm 78.00}$ $\mathbf{873.92\pm 16.03}$
Finger-spin $518.51\pm 29.03$ $\mathbf{579.27\pm 42.62}$ $\mathbf{635.18\pm 22.44}$ $\mathbf{646.37\pm 20.13}$
HalfCheetah $\mathbf{260.93\pm 47.89}$ $\mathbf{337.04\pm 48.27}$ $\mathbf{297.11\pm 29.02}$ $\mathbf{303.88\pm 38.07}$
@@
Quadruped $420.27\pm 44.00$ $\mathbf{477.29\pm 69.89}$ $\mathbf{742.05\pm 50.99}$ $\mathbf{306.78\pm 75.69}$
Table 3: Percentage improvement of RbRL-KL over RbRL across different environments and values of $n$.
Environment n=3 (%) n=4 (%) n=5 (%) n=6 (%)
Cartpole-balance $\mathbf{15.54}$ $\mathbf{3.72}$ $\mathbf{22.50}$ $\mathbf{24.37}$
Ball-in-cup $\mathbf{21.27}$ $\mathbf{9.14}$ $\mathbf{13.30}$ $\mathbf{5.47}$
Finger-spin 2.32-2.32 $\mathbf{13.24}$ $\mathbf{13.85}$ $\mathbf{15.47}$
HalfCheetah $\mathbf{60.03}$ $\mathbf{40.99}$ $\mathbf{45.24}$ $\mathbf{29.06}$
Walker $\mathbf{14.35}$ $\mathbf{22.39}$ $\mathbf{4.34}$ $\mathbf{3.42}$
Quadruped 7.51-7.51 $\mathbf{54.74}$ $\mathbf{225.98}$ $\mathbf{53.46}$
6 Limitations and Future Work
One of the limitations of the proposed approach is the variability in individual rating standards, which may be noisy due to differences in how participants interpret and evaluate the environment. Since RbRL relies on users having a basic understanding of the task and environment, these inconsistencies may reduce the reliability of the collected individual ratings. To address this, future work will involve designing a crowdsourcing framework to gather a diverse range of ratings from various participants. Developing effective noise filtering methods is important to ensure the collected data is more representative and robust, improving the reliability and accuracy of the learned models.
Although $\omega_{i}$ should be selected in a descending order with respect to the rating level, it remains an open question how to select the specific value of $\omega_{i}$. Ablation studies will be needed to optimize the performance by selecting the best $\omega_{i}$. Another interesting future research is to integrate the new KL loss terms with other RL methods in dense-reward, sparse-reward, and reward-free settings to further evaluate its effectiveness since the KL loss terms can be implemented with/without rewards. Finally, it is important to develop a systematic approach that uncovers the right number of rating classes for a given environment since a larger number of rating classes may not always lead to better performance. One potential idea is to bring the idea of Just Noticeable Difference, which quantifies the smallest change in a stimulus to be observed by humans, in psychology to determine the boundaries between different ratings. We will explore these directions as parts of our future work.